OpenStudy (anonymous):

What is 80% of 30? A. 240 B. 24 C. 2400 D. 0.24

OpenStudy (anonymous):

24

OpenStudy (anonymous):

pretty common sense

hartnn (hartnn):

a% of b is = a*b/100 so, 80% of 30 will be ?

OpenStudy (anonymous):

woah that is extremely difficult way of explaining

hartnn (hartnn):

just a different way of explaining, maybe he'll actually get it. maybe its not so obvious for him

OpenStudy (anonymous):

do .80*30

OpenStudy (anonymous):

put a decimal infront of the percent and multiply by your number

OpenStudy (anonymous):

80% = 80/100 So 80% of 30 = 30 * 80 / 100 = 24

OpenStudy (anonymous):

that seems more difficult -crumcake i see how it make 24 thank you guys :'>

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

i guess. but its less math than those ways

OpenStudy (anonymous):

it isn't always gonna be easy math so sometimes you got to learn how to work it out the hard or the long way way in my opinion

OpenStudy (anonymous):

math is always hard

OpenStudy (anonymous):

i know it is.

OpenStudy (anonymous):

just know basic math and youll be set for when youre an adult

OpenStudy (anonymous):

ok ok thank you